# Simultaneous extension of the polynomial Central Sets Theorem and the polynomial Milliken-Taylor theorem

### v1

@@ 1   Introduction
$MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)=\left\{\sum_{j=1}^{m}a_{j}\sum_{t\in F_{j}}x_{t}:F_{1}<F_{2}<\cdots<F_{m}\right\}.$
The following theorem is another version of Theorem 1.17.
@@
[2, Theorem 1.11.] Let $K\in\mathbb{N},$ let $\langle a_{j}\rangle_{j=1}^{k}$ and $\langle x_{n}\rangle_{n}$ be sequences in $\mathbb{N}.$ Let $g(z)=\sum_{j=1}^{k}a_{j}z$, and let $A\subseteq\mathbb{N}$. The following statements are equivalent.
@@
Using tensor product of ultrafilters one can find a seemingly stronger version of the Milliken-Taylor theorem.
@@
[2, Theorem 1.17] Let $(S,\cdot)$ be a semigroup, and $m\in\mathbb{N},$ and let $A\subseteq\bigtimes_{i=1}^{m}S.$ The following statements are equivalent.
@@
In [1], Beigelböck simultaneously extends the combinatorial consequence of Theorem 1.8 and Theorem 1.20. Unfortunately, Beigelböck’s result does not address the reverse direction regarding ultrafilter characterization. The following theorem is his one.
@@
Let $(S,\cdot)$ be a commutative semigroup and assume that there exists a nonprincipal minimal idempotent in $\beta S.$ For each $l\in\mathbb{N}$, let $\langle y_{l,n}\rangle_{n}$ be a sequence in $S.$ Let $k,r\in\mathbb{N}$ and let $[S]^{k}=\bigcup_{i}^{r}A_{i}.$ There exist $i\in\{1,2,\ldots,r\}$ a sequence $\langle a_{n}\rangle_{n}$ in $S$ and a sequence $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{n}<\cdots$ in $\mathcal{P}_{f}(\mathbb{N})$ such that $\Big{[}FP\big{(}\langle a_{n}\prod_{t\in\alpha_{n}}y_{g(n),t}\rangle_{n}\big{)}\Big{]}_{<}^{k}\subset A_{i}$
Our first result is to improve Theorem 1.21. We simultaneously extend Theorem 1.10 and Theorem 1.26. In our article, we restrict ourselves to the set of natural numbers, but this technique can be mimicked for commutative semigroups.
@@
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}$. Then the following two statements are equivalent.
@@
      if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{m}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
@@
    There exists a idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p.$
@@
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}$. Then the following two statements are equivalent.
@@
    There exists a idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p.$
Our next result polynomialize the Theorem 1.22. Unfortunately, we are unable to find the ultrafilter characterization of the next theorem. The main problem is that we don’t know any ultrafilter characterization of the polynomial Central Sets Theorem.
Theorem 1.24.
@@
    if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{m}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
@@
Again similar to the Corollary 1.14, the following corollary of the above theorem gives an infinitary version.
@@
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
@@
Let $m\in\mathbb{N}$, and let $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term. Then from [2, Theorem 3.2] it is clear that for every $p\in\beta\mathbb{N}$, $\tilde{h}\left(\bigotimes_{i=1}^{m}p\right)=h(p,\ldots,p).$ The following theorem is a corollary of [2, Theorem 3.2], and is the polynomial extension of Theorem 1.26.
@@
    There exists a idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p\right).$
In the above theorem, we did not consider any polynomial inside $h.$ This inspires us the following generalization of the Corollary 1.25.
@@
Let $m\in\mathbb{N}$, $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term, and let $p$ be a minimal idempotent in $(\beta\mathbb{N},+)$. Then for every $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p\right)$, and $T\in\mathcal{P}_{f}(P)$, we have two functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
@@
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P\in T$ we have
    $h\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
@@
Immediately follows from [2, Theorem 3.2], and Corollary 1.25. ∎
@@
In the next section, we prove Theorem 1.22. Let $i:\mathbb{N}\rightarrow\mathbb{N}$ be a polynomial defined by $i(x)=x$ for all $x\in\mathbb{N}.$ Then if we choose $T=\{i\}$ in Theorem 1.24, then a similar proof of Theorem 1.24 implies $(1)\implies(2)$ in Theorem 1.22. Hence we need to prove the reverse direction. So our first focus will be to prove $(2)\implies(1)$ in Theorem 1.22. Then we prove Theorem 1.24. Then we prove another version of Theorem 1.24. Then in the last section we address the polynomial extension of separating the Milliken-Taylor theorem, originally proved in [5].

### v3

@@ A Brief Review of the Algebra of the Stone-Čech Compactifications of discrete semigroups
$MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)=\left\{\sum_{j=1}^{m}a_{j}\cdot\sum_{t\in F_{j}}x_{t}:F_{1}<F_{2}<\cdots<F_{m}\right\}.$
The following theorem is another version of Theorem 1.17 that uses linear combination of idempotent ultrafilters.
@@
[2, Theorem 1.11.] Let $K\in\mathbb{N},$ let $\langle a_{j}\rangle_{j=1}^{k},$ and $\langle x_{n}\rangle_{n}$ be sequences in $\mathbb{N}.$ Let $g(z)=\sum_{j=1}^{k}a_{j}z$, and let $A\subseteq\mathbb{N}$. The following statements are equivalent.
@@
Using the tensor product of ultrafilters one can find a seemingly stronger version of the Milliken-Taylor theorem.
@@
[2, Theorem 1.17, Milliken-Taylor Theorem I] Let $(S,+)$ be a commutative semigroup, and $m\in\mathbb{N},$ and let $A\subseteq\bigtimes_{i=1}^{m}S.$ Then the following statements are equivalent.
@@
The following theorem is another version of the Milliken-Taylor theorem. Here one considers tensor products of different idempotent ultrafilters.
@@
[2, Theorem 1.16, Milliken-Taylor Theorem II] Let $(S,+)$ be a commutative semigroup, and $m\in\mathbb{N},$ and let $A\subseteq\bigtimes_{i=1}^{m}S.$ The following statements are equivalent.
  1. 1.
    For each $i\in\{1,2,\ldots,m\}$, there exist sequences $\langle x_{i,n}\rangle_{n=1}^{\infty}$ in $S$ such that
    $\left\{\left(\sum_{t\in F_{1}}x_{1,t},\sum_{t\in F_{2}}x_{2,t},\ldots,\sum_{t\in F_{k}}x_{m,t}\right):F_{1}<F_{2}<\cdots<F_{k}\right\}\subseteq A.$
  2. 2.
    There exist idempotents $p_{1},\ldots,p_{m}\in\beta S$ such that $A\in\bigotimes_{i=1}^{m}p_{i}.$
In [1], Beigelböck simultaneously extends the combinatorial consequences of both of the Theorem 1.8 and Theorem 1.20. Unfortunately, Beigelböck’s result does not address the ultrafilter characterization of sets witnessing his result. Let us first state his result, and then we address our results.
@@
Let $(S,+)$ be a commutative semigroup and assume that there exists a nonprincipal minimal idempotent in $\beta S.$ For each $l\in\mathbb{N}$, let $\langle y_{l,n}\rangle_{n}$ be a sequence in $S.$ Let $k,r\in\mathbb{N}$ and let $S]^{k}=\bigcup_{i}^{r}A_{i}.$ There exist $i\in\{1,2,\ldots,r\}$ a sequence $\langle a_{n}\rangle_{n}$ in $S$ and a sequence $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{n}<\cdots$ in $\mathcal{P}_{f}(\mathbb{N})$ such that $\Big[FS\big(\langle a_{n}\prod_{t\in\alpha_{n}}y_{g(n),t}\rangle_{n}\big)\Big]_{<}^{k}\subset A_{i}$
Our first result is the improvement of Theorem 1.22. We simultaneously extend Theorem 1.10 and Theorem 1.20 both in the algebraic and combinatorial direction. In our article, we restrict ourselves to the set of natural numbers, but this technique can be adapted for any discrete commutative semigroups.
Theorem 1.23 (Multidimensional Stronger Central Sets Theorem I).
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}^{m}$. Then the following two statements are equivalent.
@@
      if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
@@
    There exists an idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p.$
@@
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}^{m}$. Then the following two statements are equivalent.
@@
    There exists an idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p.$
Similarly to Theorem 1.21, the following theorem is a simultaneous extension of Theorem 1.10, and Theorem 1.21.
Theorem 1.25 (Multidimensional Stronger Central Sets Theorem II).
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}^{m}$. Then the following two statements are equivalent.
  1. 1.
    There exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then for every $i\in\{1,2,\ldots,m\}$, $\max H_{i}(F)<\min H_{i}(G)$, and
    2. (b)
      if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
      $\left\{\left(\sum_{i\in\beta_{1}}\left(\alpha_{1}(G_{i})+\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\cdots,\sum_{i\in\beta_{m}}\left(\alpha_{m}(G_{i})+\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right)\right\}\subset A.$
  2. 2.
    There exist idempotents $p_{1},\ldots,p_{m}$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p_{i}.$
The main difference between Theorem 1.23, and Theorem 1.25 is that in Theorem 1.23, our functions $\alpha,$ and $H$ are same throughout all the coordinates, but in Theorem 1.25, our functions $\alpha,$ and $H$ are different throughout all the coordinates. The main reason is that we used different idempotent ultrafilters. In a similar fashion to Corollary 1.24, the following corollary addresses the infinitary version of Theorem 1.25.
Corollary 1.26.
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}^{m}$. Then the following two statements are equivalent.
  1. 1.
    There exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},H_{2}\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
    2. (b)
      $\left\{\left(\sum_{i\in\beta_{1}}\left(\alpha_{1}(G_{i})+\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\cdots,\sum_{i\in\beta_{m}}\left(\alpha_{m}(G_{i})+\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right)\right\}\subset A.$
  2. 2.
    There exist idempotents $p_{1},\ldots,p_{m}$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p_{i}.$
Our next result polynomialize Theorem 1.23. Unfortunately, we are unable to find the ultrafilter characterization of the next theorem. The main problem is that we don’t know any ultrafilter characterization of the polynomial Central Sets Theorem.
Theorem 1.27 (Multidimensional Stronger Polynomial Central Sets Theorem I).
@@
    if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P_{1},\ldots,P_{m}\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
@@
The following corollary of the above theorem gives an infinitary version of Theorem 1.27. This result is the polynomial extension of Corollary 1.24.
@@
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P_{1},\ldots,P_{m}\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
@@
The following theorem is the polynomial extension of Theorem 1.25. Here we considered tensor product of different minimal idempotent ultrafilters.
Theorem 1.29 (Multidimensional Stronger Polynomial Central Sets Theorem II).
Let $m\in\mathbb{N}$, and let $p_{1},\ldots,p_{m}$ be minimal idempotents in $(\beta\mathbb{N},+)$. Then for every $A\in\bigotimes_{i=1}^{m}p_{i}$, and $T\in\mathcal{P}_{f}(P)$, we have functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H_{,}\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then for every $i\in\{1,\ldots,m\}$, we have $H_{i}(G)<H_{i}(K)$, and
  2. 2.
    if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P_{1},\ldots,P_{m}\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p_{1},\ldots,p_{m}\in E\left(\mathcal{J}_{p},+\right).$
The following corollary of the above theorem gives an infinitary version of Theorem 1.29. This result is the polynomial extension of Corollary 1.26.
Corollary 1.30.
Let $m\in\mathbb{N}$, and let $p_{1},\ldots,p_{m}$ be minimal idempotents in $(\beta\mathbb{N},+)$. Then for every $A\in\bigotimes_{i=1}^{m}p_{i}$, and $T\in\mathcal{P}_{f}(P)$, we have functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then for every $i\in\{1,2,\ldots,m\}$, we have $H_{i}(G)<H_{i}(K)$, and
  2. 2.
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P_{1},\ldots,P_{m}\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p_{1},\ldots,p_{m}\in E\left(\mathcal{J}_{p},+\right).$
@@
Let $m\in\mathbb{N}$, and let $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term. Then from [2, Theorem 3.2], it is clear that for every $p\in\beta\mathbb{N}$, $\tilde{h}\left(\bigotimes_{i=1}^{m}p\right)=h(p,\ldots,p).$ The following theorem is a corollary of [2, Theorem 3.2], and is the polynomial extension of Theorem 1.20.
@@
    There exists an idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p\right).$
In light of Corollary 1.26, the following corollary is another version of Corollary 1.32.
@@
    there exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then for each $i\in\{1,2,\ldots,m\}$, we have $\max H_{i}(F)<\min H_{i}(G)$, and
    2. (b)
      $\left\{h\left(\sum_{i\in\beta_{1}}\left(\alpha_{1}(G_{i})+\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\cdots,\sum_{i\in\beta_{m}}\left(\alpha_{m}(G_{i})+\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right)\right\}\in A.$
@@
    There exist idempotents $p_{1},\ldots,p_{m}$ in $(\mathcal{J},+)$, such that $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p_{i}\right).$
In the above theorems, we did not consider any polynomial inside $h.$ This inspires us the following generalization of the above results.
Corollary 1.34.
Let $m\in\mathbb{N}$, $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term, and let $p_{1},\ldots,p_{m}$ be minimal idempotents in $(\beta\mathbb{N},+)$. Then for every $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p_{i}\right)$, and $T\in\mathcal{P}_{f}(P)$, we have two functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then for each $i\in\{1,\ldots,m\}$, we have $H_{i}(G)<H_{i}(K)$, and
  2. 2.
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P_{1},\ldots,P_{m}\in T$ we have
    $h\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\in A.$
  3. 3.
    if for each $i\in\{1,\ldots,m\}$, $p_{i}=p$ for some minimal idempotent $p$, then we can consider $\alpha_{i}=\alpha$, and $H_{i}=H$ for some functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$.
@@
Immediately follows from [2, Theorem 3.2], and Corollaries 1.28, 1.30. ∎
@@
In the next section, we prove Theorem 1.23. Let $i:\mathbb{N}\rightarrow\mathbb{N}$ be a polynomial defined by $i(x)=x$ for all $x\in\mathbb{N}.$ Then if we choose $T=\{i\}$ in Theorem 1.27, then a similar proof of Theorem 1.27 implies $(2)\implies(1)$ in Theorem 1.23. Hence we need to prove the reverse direction. So our first focus will be to prove $(1)\implies(2)$ in Theorem 1.23. Similarly, in light of Theorem 1.29, we need to prove Theorem 1.25, we need to show only $(1)\implies(2).$ Then we prove Theorem 1.27 and its another version Theorem 1.29. In the proof of Theorem 1.27, and Theorem 1.29, we used the induction hypothesis. One can easily lift this induction hypothesis (due to the presence of idempotent ultrafilters) to conclude the Corollaries $(2)\implies(1)$ in 1.24, $(2)\implies(1)$ in 1.26, 1.28, and 1.30. However the proofs of Corollaries $(1)\implies(2)$ in 1.24, $(1)\implies(2)$ in 1.26 are similar to the proof of $(1)\implies(2)$ in Theorem 1.23, and $(1)\implies(2)$ in Theorem 1.25. Then in the last section we address the polynomial extension of separating the Milliken-Taylor theorem, originally proved in [5].